the sum of three consecutive numbers is 63.find the product of extremes.
Answer by richard1234(7193)

Let x-1, x, x+1 be the numbers. Then

(x-1) + x + (x+1) = 63

3x = 63

x = 21 --> x-1 = 20 and x+1 = 22. Their product is 440.